# SPEED: Scalable, Precise, and Efficient Concept Erasure for Diffusion Models

### v1

@@ D.2 More Baselines
Figure 8: Qualitative comparison with SPM and SPM w/o FT in erasing single and multiple instances. The erased and preserved generations are highlighted with red and green boxes, respectively. Our method demonstrates superior prior preservation compared to both SPM and SPM w/o FT. Meanwhile, without the Facilitated Transport module, SPM w/o FT shows poor erasure efficacy in multi-concept erasure (e.g., the 1st case in Snoopy, 2nd case in erasing Mickey, and 3rd case in SpongeBob) with relatively high target CS in Table 10.
D.3 Ablation Studies
Augmentation times. We ablate the augmentation times $N_{A}$ proposed in the Directed Prior Augmentation (DPA) module in Sec. 4.2, which controls the balance between rank saturation and prior coverage along with the Influence-based Prior Filtering (IPF) module. It can be observed from Fig. 9 (a) that: (1) As $N_{A}$ increases, the non-target FID exhibits a trend of first decreasing and then increasing. This suggests that when $N_{A}$ is small (i.e., $1\to 10$), augmenting existing non-target concepts with semantically similar concepts facilitates a more comprehensive prior coverage, thereby improving prior preservation. However, when $N_{A}$ exceeds a certain threshold (i.e., $10\to 20$), further augmentation of non-target concepts leads to narrowing the null-space derivation due to rank saturation, ultimately degrading prior preservation. (2) Target CS generally shows a declining trend, indicating that the proposed Prior Knowledge Refinement strategy not only improves prior preservation but also exerts a positive impact on erasure efficacy.
Augmentation ranks. Another hyperparameter to be ablated is the augmentation ranks $r$. From Eq. 6, we introduce the number of the smallest singular values, i.e., augmentation ranks $r$ in deriving $\mathbf{P}_{\text{min}}=\mathbf{U}_{\text{min}}\mathbf{U}_{\text{min}}^{\top}$ with $\mathbf{U}_{\text{min}}={\mathbf{U}}_{\mathbf{W}}[:,-r:]$. Mathematically, $r$ represents the directions in which the DPA module can augment in the concept embedding space and constrains the rank of the augmented embeddings to a maximum of $r$. As shown in Fig. 9 (b), as $r$ increases, the non-target FID exhibits an overall upward trend, indicating that introducing more ranks does not benefit prior preservation, as it narrows the null space. At the same time, as shown in Table 6, such augmentation by DPA also remains necessary, as it enables more comprehensive coverage of non-target knowledge with semantically similar concepts, leading to improved prior preservation.
Figure 9: Ablation study on two parameters, i.e., augmentation times $N_{A}$ and augmentation ranks $r$ of the DPA module. We report the target CS of erasing Van Gogh and the non-target FID averaged over other four styles (i.e., Picasso, Monet, Paul Gauguin, Caravaggio).
@@ D.3 Ablation Studies
Metirc Van Gogh Picasso Monet Paul Gauguin Caravaggio MS-COCO
CS ↓ FID ↓ FID ↓ FID ↓ FID ↓ CS ↑ FID ↓
Text Similarity 26.35 36.87 19.69 25.18 41.44 26.52 20.78
Prior Shift (Ours) 26.29 35.86 16.85 24.94 39.75 26.55 20.36
Table 12: Ablation study on the importance metrics in IPF.
@@
Importance metrics in IPF. In Sec. 4.1, we propose Importance-based Prior Filtering (IPF) in Eq. 5 and evaluate this importance with the metric prior shift $=\|\boldsymbol{\Delta}_{\text{erase}}\boldsymbol{c}\|^{2}$. Another intuitive and plausible metric is based on text similarity, e.g., the cosine similarity between each non-target embedding $\boldsymbol{c}_{0}$ and each target concept embedding $\boldsymbol{c}_{1}$, i.e., $\cos(\boldsymbol{c}_{0},\boldsymbol{c}_{1})$. Herein, we conduct an additional ablation study in terms of the metric selection in Table 12. It can be seen that text similarity can also serve as an effective metric for evaluating importance with improved non-target FID while the prior shift provides better prior preservation. This may be because text similarity is implicitly related to importance, while prior shift explicitly reflects the impact of erasure on different concepts from the model updates $\mathbf{\Delta}$. Moreover, our method can be directly scaled up to multi-concept erasure scenarios, whereas text similarity calculates $n$ similarities for $n$ target concepts, requiring additional fusion or selection strategies, introducing accumulated errors during fusion or selection.
@@ Appendix E Limitation
While SPEED demonstrates superior prior preservation, its erasure efficacy may not be as strong as some adversarial training/editing-based methods (e.g., RECE [19]), which explicitly optimize for robust concept removal. This trade-off arises from SPEED’s emphasis on maintaining non-target knowledge, potentially leading to residual traces of erased concepts in extreme cases. However, due to its efficiency and scalability in multi-concept erasure, an interesting direction for future work is to explore simultaneous erasure of adversarial examples. Given that null-space constraints inherently minimize the impact on prior knowledge, even with the addition of extra target concepts, SPEED is expected to achieve better prior preservation compared to existing methods while effectively handling adversarial concept erasure.

### v2

@@ D.2 Complete Results on Few-Concept Erasure
Figure 9: Qualitative comparison with SPM and SPM w/o FT in erasing single and multiple instances. The erased and preserved generations are highlighted with red and green boxes, respectively. Our method demonstrates superior prior preservation compared to both baselines. Meanwhile, without Facilitated Transport, SPM w/o FT shows poorer prior preservation in multi-concept erasure (e.g., marked by Refer to caption) with significant semantic changes compared to original generations.
D.3 Comparison on More Baselines
In this section, we compare against more methods because of the page limit in our main paper, including ESD [^8] https://github.com/rohitgandikota/erasing Gandikota et al. (2023), RACE [^9] https://github.com/chkimmmmm/R.A.C.E. Kim et al. (2024), Receler [^10] https://github.com/jasper0314-huang/Receler Huang et al. (2024), and SPM [^11] https://github.com/Con6924/SPM Lyu et al. (2024). While the first three methods are training-based, focusing solely on modifying model parameters, SPM not only fine-tunes the model weights using LoRA Hu et al. (2021) but also intervenes in the image generation process through Facilitated Transport. Specifically, this module dynamically adjusts the LoRA scale based on the similarity between the sampling prompt and the target concept. In other words, if the prompt contains the target concept or is highly relevant, this scale is set to a large value, whereas if there is little to no relevance, it is set close to 0, functioning similarly to a text filter. We argue that such a comparison with SPM is not fair since we only focus on modifying the model parameters, and therefore, we compare both the original SPM and SPM without Facilitated Transport (SPM w/o FT) for a fair comparison. In the latter version, the LoRA scale is set to 1 by default.
The quantitative comparative results are shown in Tables 5 and 6, where our method consistently achieves the best prior preservation compared to all compared baselines. Even equipped with Facilitated Transport (i.e., SPM w/ FT), our method achieves the lowest non-target FID (e.g., on Pikachu and Hello Kitty). This superiority amplifies as the number of target concepts increases as shown in Table 7. For example, with the number of target concepts increasing from 1 to 3, our FID in Pikachu rises from 16.81 to 21.40 (4.59 ↑), while SPM increases from 19.82 to 37.51 (17.69 ↑), where a similar pattern is observed in Hello Kitty (Our 4.48 ↑ v.s. SPM’s 15.68 ↑). Once removing the Facilitated Transport module, SPM w/o FT shows poorer prior preservation with rapidly increasing FIDs (highlighted in red in Table 5). More qualitative results are shown in Fig. 9.
D.4 On Implicit Concept Erasure
Table 8: Evaluation of implicit concept erasure in erasing nudity on four benchmarks. We report the Attack Success Rate (ASR) detected by NudeNet with a threshold of 0.6. ✓ and $\times$ indicate whether the method can defend against white-box attacks, respectively. The best and second-best results are marked in bold and underlined.
I2P MMA Ring-A-Bell UnlearnDiff Time (s) ↓ MS-COCO White-Box Attack
CS FID
MACE Lu et al. (2024a) 0.21 0.04 0.05 0.67 55 (×15) 24.06 52.78
CPE Lee et al. (2025b) 0.07 0.01 0.00 - 500 (×138) 26.32 48.23 $\times$
AdvUnlearn Zhang et al. (2024b) 0.04 0.00 0.00 0.21 15860 (×4400) 24.05 57.22
UCE Gandikota et al. (2024) 0.24 0.38 0.39 0.80 1.2 (×0.33) 26.24 38.60
RECE Gong et al. (2025) 0.14 0.20 0.18 0.65 1.5 (×0.41) 25.98 40.37
RACE Kim et al. (2024) 0.23 0.29 0.21 0.47 2910 (×800) 25.54 42.73
Receler Huang et al. (2024) 0.13 0.07 0.01 - 5560 (×1500) 25.93 40.29 $\times$
Ours w/o AT 0.20 0.24 0.20 0.75 3.6 (×1) 26.29 37.82
Ours w/ AT 0.10 0.01 0.00 0.45 4.5 (×1.25) 26.03 39.51
Evaluation setup. We evaluate the erasure efficacy on implicit concepts (e.g., nudity), where the target concept does not explicitly appear in the text prompt. We conduct experiments on the Inappropriate Image Prompt (I2P) benchmark Schramowski et al. (2023), which consists of various implicit inappropriate prompts involving violence, sexual content, and nudity. To evaluate adversarial robustness, we further introduce three adversarial attack benchmarks, including two black-box benchmarks (MMA Yang et al. (2024b) and Ring-A-Bell Tsai et al. (2023)) and one white-box benchmark (UnlearnDiff Zhang et al. (2024c)). For concept erasure, we follow the setting in Gong et al. (2025) to erase nudity $\to$ ‘ ’. During evaluation, we use NudeNet Bedapudi (2019) with a threshold of 0.6 to detect nude content and report the Attack Success Rate (ASR).
Analysis and discussion. In addition to the aforementioned methods, we introduce additional adversarial training-based methods for comparison, including CPE Lee et al. (2025b), AdvUnlearn Zhang et al. (2024b), RACE Kim et al. (2024), and Receler Huang et al. (2024). These methods enhance robustness against adversarial attacks by explicitly incorporating adversarial training objectives. We also adapt our method with adversarial training/editing (denoted as Ours w/ AT) following the setting in RECE Gong et al. (2025) to provide a fair comparison. As shown in Table 8, we observe that adversarial training-based methods such as CPE and AdvUnlearn achieve strong erasure efficacy but incur extremely high computational costs. Editing-based approaches like UCE and RECE are more efficient yet less robust under both black-box and white-box attacks. Notably, CPE and Receler rely on additional modules for concept erasure, which makes them particularly vulnerable in the white-box setting since attackers can directly exploit these components to bypass erasure. In contrast, our method without adversarial training (Ours w/o AT) already offers a favorable balance between efficiency and prior preservation, and extending it with adversarial training/editing (Ours w/ AT) further improves robustness, reducing ASR across all benchmarks and lowering the white-box UnlearnDiff score from 0.75 to 0.45 while maintaining competitive runtime and prior knowledge preservation.
D.5 Ablation Studies
@@
((a))
((b))
((c))
Figure 10: Ablation study on hyperparameters. We report target CS of erasing Van Gogh and non-target FID averaged over other four styles (i.e., Picasso, Monet, Paul Gauguin, Caravaggio).
Filtering scale. We ablate the filtering threshold scale in the Influence-based Prior Filtering (IPF) module in Sec. 4.1 by scaling the impact scores $\mu$ with a factor $\alpha$, which controls the strength of filtering influential priors. As shown in Fig. 10 (a), varying $\alpha$ directly affects the trade-off between erasure efficacy and prior preservation. When $\alpha$ is small (i.e., close to 0), more weakly affected priors are included in the retain set, increasing its rank and overly shrinking the null space. This leads to worse erasure efficacy (higher CS) and poor preservation (higher FID). Conversely, a higher $\alpha$ yields better erasure performance due to fewer retain concepts, but still increases the FID because of non-comprehensive prior coverage. The best balance is observed at moderate thresholds (e.g., $\alpha=1$ in our setup), achieving both effective erasure and competitive prior preservation.
Augmentation times. We ablate the augmentation times $N_{A}$ proposed in the Directed Prior Augmentation (DPA) module in Sec. 4.2, which controls the balance between semantic degradation and retain coverage along with the Influence-based Prior Filtering (IPF) module. It can be observed from Fig. 10 (b) that: (1) As $N_{A}$ increases, the non-target FID exhibits a trend of first decreasing and then increasing. This suggests that when $N_{A}$ is small (i.e., $1\to 10$), augmenting existing non-target concepts with semantically similar concepts facilitates a more comprehensive retain coverage, thereby improving prior preservation. However, when $N_{A}$ exceeds a certain threshold (i.e., $10\to 20$), further augmentation of non-target concepts leads to narrowing the null-space derivation with semantic degradation, ultimately degrading prior preservation. (2) Target CS generally shows a declining trend, indicating that the proposed Prior Knowledge Refinement strategy not only improves prior preservation but also exerts a positive impact on erasure efficacy.
Augmentation ranks. Another hyperparameter to be ablated is the augmentation ranks $r$. From Eq. 7, we introduce the number of the smallest singular values, i.e., augmentation ranks $r$ in deriving $\mathbf{P}_{\text{min}}=\mathbf{U}_{\text{min}}\mathbf{U}_{\text{min}}^{\top}$ with $\mathbf{U}_{\text{min}}={\mathbf{U}}_{\mathbf{W}}[:,-r:]$. Mathematically, $r$ represents the directions in which the DPA module can augment in the concept embedding space and constrains the rank of the augmented embeddings to a maximum of $r$. As shown in Fig. 10 (c), as $r$ increases, the non-target FID exhibits an overall upward trend, indicating that introducing more ranks does not benefit prior preservation, as it narrows the null space. At the same time, as shown in Table 3, such augmentation by DPA also remains necessary, as it enables more comprehensive coverage of non-target knowledge with semantically similar concepts, leading to improved prior preservation.
@@ Appendix E Limitation
Despite the promising results, our work still has limitations. SPEED relies on linear null-space projections, which may not fully capture the nonlinear interactions between concepts in large diffusion models. In practice, this can lead to imperfect preservation when erasing highly entangled or stylistically subtle concepts. In addition, our evaluation mainly covers benchmarks with explicit or moderately implicit concepts; the effectiveness on more abstract (e.g., freedom), compositional (e.g., a blue cat), or culturally nuanced (e.g., Día de los Muertos) concepts remains less explored. Finally, although our method scales efficiently to 100 concepts, extending it to even larger-scale or continual erasure settings may require additional mechanisms.